0.220 grams of chromium was burned in air and reacted with the oxygen to give 0.322 grams of chromium oxide. What is the empirical formula of the oxide? What is the name of the oxide?
Please include steps.
1
Expert's answer
2020-09-14T08:22:51-0400

QUESTION # 133018

0.220 grams of chromium was burned in air and reacted with the oxygen to give 0.322 grams of chromium oxide. What is the empirical formula of the oxide? What is the name of the oxide?

ANSWER

Mass of chromium burned in air = 0.220g

Mass of oxygen used = 0.322-0.220= 0.102g

Element                               Cr                                           O

Mass                                      0.220                                      0.102

Atomic mass                         51.9961                                  15.999

Moles of each element         0.220/51.9961=0.004231      0.102/15.999=0.006375

Mole ratio                            0.004231/0.004231=1           0.006375/0.00423=1.5          

Whole number ratio            1×2=2                                     1.5×2=3

Empirical formula is Cr2O3    

The name of oxide is Chromium (III) oxide